# Low temperature decoherence dynamics in molecular spin systems using the Lindblad master equation

### v1

@@ I Introduction
Molecular spin systems are centerpieces of a variety of emerging technologies including quantum computing, sensing, single-molecule magnets, and spintronics. [1, 2, 3, 4, 5, 6, 7, 8] In addition to technological applications, molecular spin systems are fundamental components of contemporary problems in experimental and theoretical chemistry across many fields. In cold chemistry, for example, the nuclear spin states of ultracold molecular collisions can influence the branching ratios of product channels. [9, 10, 11, 12] Other recent experiments characterize the spin dynamics in molecular systems by measuring the spin-lattice, $T_{1}$, and spin-spin, $T_{2}$, relaxation rates. [13, 14, 15, 16, 17, 18, 19, 20] In the context of quantum computing magnetic molecules could serve as quantum bits, or qubits, due to their optical addressability, synthetic tunability, and potential for high-temperature operation. [21, 22, 23, 24, 25] Fundamental understanding of the electron dynamics and the manipulation and control of the electronic state are essential for effective use of molecular spin systems in practical settings.
@@
In this Article, we develop a phenomenological model capable of describing ensemble dynamics of molecular spin systems interacting with an environment. Because these systems undergo relaxation, one natural approach is to use an open quantum system perspective. We focus on the low-temperature setting using the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation to simulate the quantum dynamics. [27, 28, 29, 30] We utilize ab initio electronic structure to incorporate the geometric and dynamic distributions present in ensemble systems to predict $T_{2}$ relaxation. We demonstrate the utility of this approach by simulating the decoherence dynamics of two molecular spin systems. This approach is general to cold molecular spin systems undergoing decoherence dynamics with applications in a wide array of emerging quantum technologies.
@@ II Theory and Methods
In either a time- or ensemble-averaged experiment, this time-dependent magnetic field rescales the Zeeman energy of the electron, broadening the energy states by dynamically changing the precession frequency of the electron. In a spin echo experiment at low temperatures, decoherence due to static inhomogeneities of an external magnetic field is removed, but the dynamic flip-flop broadening processes will remain. The time-dependent field induced by the nuclear spin flip-flops is therefore the most significant remaining contributor to dephasing. [34, 35, 36] Modeling the ensemble spin-decoherence in these systems is therefore challenging due to the influence of the static electronic structure along with the irreversible nature of the electron broadening process.
Here, we use the GKSL equation enhanced with ab initio electronic structure to model the spin decoherence of an ensemble of molecules,
@@
All electronic structure calculations were carried out using the ORCA 5 quantum chemistry software package. [37] Since the vanadium-oxo complexes are not strongly correlated, density functional theory (DFT) is sufficient to describe their electronic structure. The B3LYP hybrid functional was used, in addition to a triple-$\zeta$ polarized basis set for the carbon, hydrogen, oxygen, and sulfur, with a core-polarized CP(PPP) basis set for the vanadium ion, in addition to an auxiliary basis set. [38, 39, 40, 41] For each molecule V1-V4 in Figure 2, geometries were optimized, and from those equilibrium geometries, vibrational modes were calculated. From the trajectory of these normal modes, two geometries were selected for each mode in each molecule, and single-point calculations using the above DFT parameters were performed. During each calculation, hyperfine tensors were computed, taking into account dipole-dipole interactions, spin-orbit coupling, and the Fermi contact interaction between the electron and the spin-active hydrogen and vanadium nuclei in the molecules. [32, 33, 31] The same process was completed for CuS and CuSe, using the M06L functional with an ANO-RCC-DZP basis for the carbon and hydrogen, and an ANO-RCC-TZP basis for the copper, sulfur, and selenium, in addition to auxiliary basis sets. [42, 43, 44, 45, 46, 47]
From the atomic coordinates of each geometry, the rates of flip-flops for each nuclear pair $i,j$ in a given ensemble was calculated according to,
@@
where the $n$ index runs over all the nuclei in the molecule that are not part of the current flip-flop pair $ij$[48] $I$ is the spin of the bath nuclei, which is $\frac{1}{2}$ in the present work. For this work, vanadium and copper are approximated as a spin $\frac{1}{2}$ nucleus. This is not expected to be a significant source of error for the present model, as the difference in hyperfine energy between the vanadium and distal hydrogens is so large that nuclear flip-flops between the vanadium and other spin-active nuclei are prevented, and therefore do not contribute to decoherence of the electron. This equation is derived by considering the lineshape of a nuclear pair in a nuclear magnetic resonance (NMR) experiment. $\kappa_{ij}$ represents broadening of the NMR lineshape of a nuclear pair because of the dipolar interactions between the nuclear flip-flop pair and all other spin-active nuclei. The broadening of the lineshape allows the flip-flop process to occur even when the Zeeman energy of each nuclei differs. [49, 50, 51, 52, 53] The terms $A_{i}$ and $A_{j}$ represent the center of each feature in the NMR lineshape, or the chemical shift, of the two nuclei, and are taken from the hyperfine tensors calculated with the electronic structure methods described above. Data was generated by setting $\Delta$=$\Delta_{0}$=00 and $\Delta$=$\Delta_{ij}$=$|A_{i}-A_{j}|$ in Eq. 2. For each pair $ij$ the distribution of flip-flop rates is given by the set $\{T_{ij}\}$, from which the standard deviation of flip-flop rates, $\sigma_{\{T_{ij}\}}$ for each pair in an ensemble is calculated.
@@
To treat the Liouvillian part of the dynamics, some approximation is necessary. The vanadium-oxo molecular systems contain 13 spin-active nuclei, in addition to the electron, which amounts to a $2^{14}$ dimensional Hilbert space, while the copper systems contain 9 spin-active nuclei in addition to the electron, for a $2^{10}$ dimensional Hilbert space. To minimize this cost, the Hamiltonian is factorized into clusters of nuclear spins in their interactions with the central electron, following the cluster-correlation expansion (CCE). [55, 56, 57, 58, 59] For the purposes of this work, clusters of two nuclear spins are sufficient since the mechanism of dephasing being considered here is the pairwise flip-flop of two nuclear spins at a time. [58] Finally, the Hamiltonian parameters that were factorized using CCE were taken from the equilibrium geometry of each molecule calculated using DFT.
@@ III Results
In Figure 3 we show the calculated coherence profiles using the GKSL equation for each ensemble of V1-V4, with the experimental data shown in the inset. [15] The parameters given by fits to these coherence decays are shown in Table 1.
Figure 3: Simulated decoherence of the vanadium-oxo series using the GKSL equation, with experimental data reproduced from Ref. 15 shown in the inset.
@@
To further evaluate the impact of $\Delta$ on the spread of flip-flop rates, $\sigma_{\{T_{ij}\}}$ is plotted in Figure 5 with $\Delta_{ij}$ and $\Delta_{0}$ for the nuclear pairs that have $\sigma_{\{T_{ij}\}}$ $>$ 5 kHz.
@@
The example of the vanadium ensembles demonstrates the importance of the location of the spin-active nuclei relative to the electron in determining decoherence. Altering the local magnetic environment of the electron can also be achieved by changing the ligand that is bonded to the metal center. To demonstrate this we compare the coherence profiles of a copper molecular spin system with sulfur- or selenium-based ligands shown in Figure 6, labelled as CuS and CuSe, respectively. Experiments show that CuSe has a longer-lived coherence due to increased copper-ligand covalency, indicating that the electron is delocalized on to the selenium ligand. [43]
@@
Table 2: Decay parameters ($ms$) for CuS and CuSe, along with experimental decay constants from Ref. 43.
@@ IV Discussion and Conclusions
Through these two molecular spin systems, we have shown that the GKSL equation qualitatively reproduces experimental trends in spin decoherence. Importantly, our model includes information about the local magnetic environment of the decohering spin through electronic structure calculations. The magnetic environment can be tuned in various ways, for example through increased metal-ligand covalency or through the internuclear distance of the spins. As an example of the former, we compare two copper molecules to show how increasing metal-ligand covalency can impact spin-spin decoherence. Increased metal-ligand covalency, or delocalization of the electron onto the selenium ligands, increases the hyperfine coupling via the Fermi contact and dipole-dipole interactions between the electron and the nuclear spins in the copper molecules. As an example of the latter, we compare a series of vanadium-oxo complexes showing the smallest complex exhibited the longest coherence time. Our results validate conventional understanding of decoherence in these molecular species; namely, that a spin diffusion barrier mitigates decoherence processes when nuclear spins are near the electronic spin. Our approach naturally incorporates the distinct magnetic environments of each nuclei by using ab initio electronic structure in its treatment of the nuclear pair flip-flop process. In doing so, the presence of the spin diffusion barrier is a natural outcome, providing a theoretical justification for its use in explaining observed molecular $T_{2}$ trends.
@@
V Supporting Information
Supporting Information includes the full standard deviation of the flip-flop rate for each nuclear pair in the ensembles V1-V4. The fitting function information used to extract decay times is also shown.
VI Acknowledgements
K.H.M. acknowledges start-up funding from Washington University in St. Louis and the University of Minnesota.

### v2

@@ I Introduction
Molecular spin systems are centerpieces of a variety of emerging technologies including quantum computing, sensing, single-molecule magnets, and spintronics. [1, 2, 3, 4, 5, 6, 7, 8, 9] In addition to technological applications, molecular spin systems are fundamental components of contemporary problems in experimental and theoretical chemistry across many fields. In cold chemistry, for example, the nuclear spin states of ultracold molecular collisions can influence the branching ratios of product channels. [10, 11, 12, 13] Other recent experiments characterize the spin dynamics in molecular systems by measuring the spin-lattice, $T_{1}$, and spin-spin, $T_{2}$, relaxation rates. [14, 15, 16, 17, 18, 19, 20, 21] In the context of quantum computing magnetic molecules could serve as quantum bits, or qubits, due to their optical addressability, synthetic tunability, and potential for high-temperature operation. [22, 23, 24, 25, 26] Fundamental understanding of the electron dynamics and the manipulation and control of the electronic state are essential for effective use of molecular spin systems in practical settings.
@@
In this Article, we develop a phenomenological model capable of describing ensemble dynamics of molecular spin systems interacting with an environment. Because these systems undergo relaxation, one natural approach is to use an open quantum system perspective. We focus on the low-temperature setting using the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation to simulate the quantum dynamics. [28, 29, 30, 31] We incorporate the geometric and dynamic distributions present in ensemble systems to predict $T_{2}$ relaxation. Specifically, we use ab initio electronic structure to determine the GKSL decoherence rates for non-interacting ensembles of molecular spins. We demonstrate the utility of this approach by simulating the decoherence dynamics of two molecular spin systems. This approach is general to cold molecular spin systems undergoing decoherence dynamics with applications in a wide array of emerging quantum technologies.
@@ II Theory and Methods
In either a time- or ensemble-averaged experiment, this time-dependent magnetic field rescales the Zeeman energy of the electron, broadening the energy states by dynamically changing the precession frequency of the electron. In a spin echo experiment at low temperatures, decoherence due to static inhomogeneities of an external magnetic field is removed, but the dynamic flip-flop broadening processes will remain. The time-dependent field induced by the nuclear spin flip-flops is therefore the most significant remaining contributor to dephasing. [35, 36, 37, 38] Modeling the ensemble spin-decoherence in these systems is therefore challenging due to the influence of the static electronic structure along with the irreversible nature of the electron broadening process.
Here, we use the GKSL equation enhanced with ab initio electronic structure to model the spin decoherence of an ensemble of molecules. The GKSL equation is a commonly used open quantum system master equation derived under the Born-Markov approximation, meaning it can be used to capture the dissipative dynamics in quantum systems weakly coupled to quickly relaxing environments. [29, 30, 28, 31] This equation has found application in many fields including quantum biology, quantum information, quantum optics, and condensed matter physics. [39, 40, 41, 42, 43, 44, 31] A major benefit of the GKSL equation is that, by construction, it preserves the positivity of the density matrix. [29, 30, 28, 31] This equation takes the form,
@@
In typical ensemble experiments measurements will sample a distribution of molecular configurations or geometries. For non-interacting systems at low temperature, the distribution of geometries can be approximated from the zero-point vibrational modes of a molecule. To simulate such an ensemble, we sample the zero-point vibrational modes of each molecule using density functional theory (DFT), which results in a distribution of local magnetic environments for the nuclear spins. Computing this distribution is an important part of our approach because it estimates the broadening of the rescaled Zeeman energy in Fig. 1 which is the source of the irreversible decoherence in a molecular ensemble. For each molecule we generate $2\cdot(3N-6)$ different geometries from the normal modes.
We carry out all electronic structure calculations using the ORCA 5 quantum chemistry software package. [45] Since the vanadium-oxo complexes are not strongly correlated, DFT is sufficient to describe their electronic structure. We used the B3LYP hybrid functional, in addition to a triple-$\zeta$ polarized basis set for the carbon, hydrogen, oxygen, and sulfur, with a core-polarized CP(PPP) basis set for the vanadium ion, in addition to an auxiliary basis set. [46, 47, 48] For each molecule V1-V4 in Figure 2, we optimized geometries optimized, and from those equilibrium geometries, we calculated vibrational modes calculated. We followed the same process for CuS and CuSe, using the M06L functional with an ANO-RCC-DZP basis for the carbon and hydrogen, and an ANO-RCC-TZP basis for the copper, sulfur, and selenium, in addition to auxiliary basis sets. [49, 50, 51, 52, 53, 54] The functionals and basis sets were chosen following previous work which qualitatively matches experimentally measured spin densities. [55, 56]
During each calculation, we computed hyperfine tensors, taking into account dipole-dipole interactions, spin-orbit coupling, and the Fermi contact interaction between the electron and the spin-active hydrogen and vanadium nuclei in the molecules. We computed the hyperfine tensors via an expectation value taken over the spin density of each molecule. [45, 33, 34, 32] We used the resulting hyperfine tensors to compute the difference in magnetic environment between two nuclear spins $i$ and $j$ as $\Delta_{ij}=|A_{i}-A_{j}|$, with full details for electronic structure methods and computation of hyperfine tensors shown in the Supplementary Information (SI).
Following an analogous process for nuclear spin lattices with defects, [57] we calculate the flip-flop rate for each nuclear pair $i,j$ in a specific geometry according to,
@@
where the $n$ index runs over all the nuclei in the molecule that are not part of the current flip-flop pair $ij$[57] $I$ is the spin of the bath nuclei, which is $\frac{1}{2}$ in the present work. For this work, vanadium and copper are approximated as a spin $\frac{1}{2}$ nucleus. This is not expected to be a significant source of error for the present model, as the difference in hyperfine energy between the vanadium and copper metal ions and the distal hydrogens is so large that nuclear flip-flops between the vanadium and other spin-active nuclei are prevented, and therefore do not contribute to decoherence of the electron.
This equation is derived by considering the lineshape of a nuclear pair in a nuclear magnetic resonance (NMR) experiment. [57] $\kappa_{ij}$ represents broadening of the NMR lineshape of a nuclear pair because of the dipolar interactions between the nuclear flip-flop pair and all other spin-active nuclei. The broadening of the lineshape allows the flip-flop process to occur even when the Zeeman energy of each nuclei differs. [58, 59, 60, 61, 62] The terms $A_{i}$ and $A_{j}$ represent the center of each feature in the NMR lineshape, or the chemical shift, of the two nuclei, and are taken from the hyperfine tensors calculated with the electronic structure methods described above. Flip-flop rates were generated by setting $\Delta$=$\Delta_{0}$=00 and $\Delta$=$\Delta_{ij}$=$|A_{i}-A_{j}|$ in Eq. 2. For each pair $ij$ the distribution of flip-flop rates present in an ensemble solution is given by the set $\{T_{ij}\}$, from which the standard deviation of flip-flop rates, $\sigma_{\{T_{ij}\}}$ for each pair in an ensemble is calculated. Calculating the distribution of flip-flop rates and their use as the decay parameter for molecular spin decoherence is a major contribution of this work. This provides a relationship between ab initio quantum chemistry and the rate of nuclear spin flip-flops in molecular spin systems.
@@
To treat the Liouvillian part of the dynamics, some approximation is necessary. The vanadium-oxo molecular systems contain 13 spin-active nuclei, in addition to the electron, which amounts to a $2^{14}$ dimensional Hilbert space, while the copper systems contain 9 spin-active nuclei in addition to the electron, for a $2^{10}$ dimensional Hilbert space. To minimize this cost, the Hamiltonian is factorized into clusters of nuclear spins in their interactions with the central electron, following the master equation cluster-correlation expansion (ME-CCE). [64, 65, 66, 67, 68, 69, 70] A major contribution of this work is the physically-motivated derivation of the decay rates $\gamma_{K}$, where previous approaches as well as implementations of ME-CCE have used $\gamma_{K}$ as an adjustable parameter to match experiment. [69, 38] Here, no adjustable parameters are present, and qualitative experimental agreement is achieved. For the purposes of this work, clusters of two nuclear spins are sufficient since the mechanism of dephasing being considered here is the pairwise flip-flop of two nuclear spins at a time. [68] Finally, the Hamiltonian parameters that were factorized using CCE were taken from the equilibrium geometry of each molecule calculated using DFT. Additional details on the calculation of spin dynamics is included in the SI.
@@ III Results
In Figure 3 we show the calculated coherence profiles using the GKSL equation for each ensemble of V1-V4, with the experimental data shown in the inset. [16] Experimental decoherence profiles are generally fit to an empirical equation (see SI) which provides an estimate of the decay constant. Following this procedure, Table 1 shows the resulting decay constants which are fit from our simulated data.
Figure 3: Simulated decoherence of the vanadium-oxo series using the GKSL equation, with experimental data reproduced from Ref. 16 shown in the inset.
@@
To further evaluate the impact of $\Delta$ on the spread of flip-flop rates, $\sigma_{\{T_{ij}\}}$ is plotted in Figure 5 with $\Delta_{ij}$ and $\Delta_{0}$ for the nuclear pairs that have $\sigma_{\{T_{ij}\}}$ $>$ 5 kHz. $\sigma_{\{T_{ij}\}}$ for all nuclear spins are shown in the SI.
@@
The example of the vanadium ensembles demonstrates the importance of the location of the spin-active nuclei relative to the electron in determining decoherence. Altering the local magnetic environment of the electron can also be achieved by changing the ligand that is bonded to the metal center. To demonstrate this we compare the coherence profiles of a copper molecular spin system with sulfur- or selenium-based ligands shown in Figure 6, labelled as CuS and CuSe, respectively. Experiments show that CuSe has a longer-lived coherence due to increased copper-ligand covalency, indicating that the electron is delocalized on to the selenium ligand. [55] Computed spin densities for each molecule using the previously described DFT parameters are shown in the SI.
@@
Table 2: Decay parameters ($ms$) for CuS and CuSe, along with experimental decay constants from Ref. 55.
@@ IV Discussion and Conclusions
Through these two sets of molecular spin systems, we have shown that the GKSL equation qualitatively reproduces experimental trends in spin decoherence. Importantly, the local magnetic environment of the decohering spin through electronic structure calculations which directly relates to the decoherence rate of molecular spin ensembles. The magnetic environment can be tuned in various ways, for example through increased metal-ligand covalency or through the internuclear distance of the spins. As an example of the former, we compare two copper molecules to show how increasing metal-ligand covalency can impact spin-spin decoherence. Increased metal-ligand covalency, or delocalization of the electron onto the selenium ligands, increases the hyperfine coupling via the Fermi contact and dipole-dipole interactions between the electron and the nuclear spins in the copper molecules. As an example of the latter, we compare a series of vanadium-oxo complexes showing the smallest complex exhibited the longest coherence time. Our results validate conventional understanding of decoherence in these molecular species; namely, that a spin diffusion barrier mitigates decoherence processes when nuclear spins are near the electronic spin. Our approach naturally incorporates the distinct magnetic environments of each nuclei by using ab initio electronic structure in its treatment of the nuclear pair flip-flop process. In doing so, the presence of the spin diffusion barrier is a natural outcome, providing a theoretical justification for its use in explaining observed molecular $T_{2}$ trends.
@@
V Supplementary Information
Supplementary Information includes the full standard deviation of the flip-flop rate for each nuclear pair in the ensembles V1-V4. The fitting function information used to extract decay times from simulated data is also provided. Additional details on the electronic structure methods used and the ME-CCE method used are provided, including the computed spin densities of each molecule, calculation of hyperfine tensors, and the effective spin Hamiltonian used to calculate spin dynamics.
VI Author Contributions
CRediT: Timothy J. Krogmeier:data curation, formal analysis, methodology, validation, visualization, writing-original draft, writing-review & editing; Anthony W. Schlimgen: conceptualization, formal analysis, methodology, software, supervision, writing-original draft, writing-review & editing; Kade Head-Marsden: conceptualization, formal analysis, funding acquisition, methodology, project administration, resources, software, supervision, writing-original draft, writing-review & editing.
VII Data Availability Statement
Additional data and code that support the findings of this study are available from the corresponding author upon reasonable request and on the Head-Marsden group’s GitHub respository.